# AirSLAM: An Efficient and Illumination-Robust Point-Line Visual SLAM System

### v2

@@ VII-C2 Dynamic Illumination
VII-C3 Low Illumination
$V_{out}=AV_{in}^{\frac{1}{\gamma}},$ (22)
We present the comparison result in Fig. 10. As the errors of PL-SLAM are much greater than other methods, we do not show its result. Tracking failures and large drift errors, i.e., the RMSE is more than 1 m , are also marked. It can be seen that low illumination has varying degrees of impact on different systems. Our system and Basalt achieve the best outcome. Basalt is more stable for low illumination: its RMSE remains almost unchanged under different brightness levels. Our system is more accurate: AirSLAM has the smallest error on most sequences. The RMSEs of ORB-SLAM3 and OKVIS increase as the brightness decreases. They even experience tracking failures or large drift errors on $L_{10}$ and $L_{11}$.
@@ VII-D3 Results
TABLE VI: Ablation study. The recall rates (%) of our system with and without the structure graph in the map reuse module.
@@
VII-D4 Analysis
We find that the VPR methods perform very well on the first type of image but perform poorly on the second type of image. Therefore, their recall rates are very low on P001 and P006, which contain more of the second type of images. This may be because their global features are usually trained on datasets that have a lot of distinct buildings and landmarks, which makes them rely more on such semantic cues to retrieve similar images. By contrast, our system is based on the DBoW method, which only utilizes the low-level local features of images, so it achieves similar performance on the first and second types of images. This also proves the good generalization ability of our system. However, neither our system nor VPR methods can process the pairs with little overlap due to relying on local feature matching. Such image pairs are abundant in the P011.
@@ VII-E Ablation Study
In this part, we verify the effectiveness of the relocalization method. This experiment is conducted on the TartanAir Day/Night Localization dataset. We compare the systems with and without the second step proposed in Section VI-B2. The results are presented in Table VI, where w/o G. denotes our system without the structure graph, and $N_{\mathcal{C}}$ denotes the candidate number for local feature matching. It shows that using junctions, line features, and structure graphs to filter out relocalization candidates significantly improves recall rates. AirSLAM outperforms w/o G. across all sequences, and when $N_{\mathcal{C}}$ is 3, 5, and 10, the average improvements are 9.6%, 6.0%, and 3.8%, respectively, which demonstrates the effective performance of the proposed method.
@@
Figure 13: For odometry efficiency, we compare with the traditional methods and disable the loop detection from all the methods. For mapping efficiency, we compute the total mapping time and compare it with Hloc.
VII-F Efficiency Analysis
Efficiency is essential for robotics applications, so we also evaluate the efficiency of the proposed system. We first compare the running time of our AirSLAM with several SOTA VO and SLAM systems on a computer with an Intel i9-13900 CPU and an NVIDIA RTX 4070 GPU. Then we deploy AirSLAM on an NVIDIA Jetson AGX Orin to verify the efficiency and performance of our system on the embedded platform.
VII-F1 Odometry Efficiency
The VO/VIO efficiency experiment is conducted on the MH_01_easy sequence of the EuRoC dataset. We compare our AirSLAM with several SOTA systems. The loop detection and GBA are disabled from all the systems for a fair comparison. The metrics are the runtime per frame and the CPU usage. The results are presented in Fig. 13a, where 100% CPU usage means 1 CPU core is utilized. It should be additionally noted that DROID-SLAM actually uses 32 CPU cores, and its CPU usage in Fig. 13a is only for a compact presentation. Our system is the fastest among these systems, achieving a rate of 73 FPS. In addition, due to extracting and matching features using the GPU, our system requires relatively less CPU resources. We also test the GPU usage. It shows that DROID-SLAM requires about 8GB GPU memory, while our AirSLAM only requires around 3GB.
@@ VII-F3 Embedded Platform
TABLE VII: Efficiency comparison of our system on two platforms.
@@
We use 8 sequences in the EuRoC dataset to evaluate the efficiency of AirSLAM on the embedded platform. The suffixes, i.e., “-Jetson” and “-PC”, are added to distinguish results on different platforms. On the Jetson, we modify three parameters in our system to improve efficiency. First, we reduced the number of detected keypoints from 350 to 300. Second, we change two parameters in Section V-D to make keyframes sparser, i.e., $\alpha_{1}$ and $\alpha_{2}$ are changed from 0.65 and 0.1 to 0.5 and 0.2, respectively. The other parameters are the same on these two platforms. The comparisons of efficiency and absolute trajectory error (ATE) are presented in Table VII and Fig. 14, respectively. Our AirSLAM can run at a rate of $40\mathrm{Hz}$ on the Jetson while only consuming 2 CPU cores and 989MB GPU memory. We find the runtime of the offline map optimization is very close on these two platforms. This is because AirSLAM-Jetson selects fewer keyframes than AirSLAM-PC, so the loop closure and GBA are faster.
@@ VIII Conclusion
In this work, we present an efficient and illumination-robust hybrid vSLAM system. To be robust to challenging illumination, the proposed system employs the CNN to detect both keypoints and structural lines. Then these two features are associated and tracked using a GNN. To make the system more efficient, we propose PLNet, which is the first unified model to detect both point and line features simultaneously. Furthermore, a multi-stage relocalization method based on both appearance and geometry information is proposed for efficient map reuse. We design the system with an architecture that includes online mapping, offline optimization, and online relocalization, making it easier to deploy on real robots. Extensive experiments show that the proposed system outperforms other SOTA vSLAM systems in terms of accuracy, efficiency, and robustness in illumination-challenging environments.
Despite its remarkable performance, the proposed system still has limitations. Like other point-line-based SLAM systems, our AirSLAM relies on enough line features, so it is best to apply it to man-made environments. This is because our system was originally designed for warehouse robots. In the unstructured environments, it will degrade into a point-only system.

### v4

@@ VII-C Mapping Robustness
VII-C1 Onboard Illumination
VII-C2 Dynamic Illumination
The UMA-VI dataset is a visual-inertial dataset gathered in challenging scenarios with handheld custom sensors. We selected sequences with illumination changes to evaluate our system. As shown in Fig. 9, it contains many subsequences where the image suddenly goes dark due to turning off the lights. In the subsequence of the top row in Fig. 9, the total darkness lasted for 0.16 s , and low illumination lasted for 0.64 s . In the subsequence of the bottom row, the total darkness lasted for 0.32 s , and low illumination lasted for 0.4 s . It is more challenging than the OIVIO dataset for vSLAM systems. As the ground-truth poses are only available at the beginning and the end of each sequence, we disabled the loop closure part from all the evaluated methods.
VII-C3 Low Illumination
$V_{out}=AV_{in}^{\frac{1}{\gamma}},$ (22)
We present the comparison result in Fig. 11. As the errors of PL-SLAM are much greater than other methods, we do not show its result. Tracking failures and large drift errors, i.e., the RMSE is more than 1 m , are also marked. It can be seen that low illumination has varying degrees of impact on different systems. In this scenario, optical flow-based methods demonstrate greater stability compared to feature-based methods. Nonetheless, our approach achieves comparable performance to Basalt and DROID-SLAM, which utilize sparse and dense optical flow, respectively. The RMSEs of ORB-SLAM3 and OKVIS increase as the brightness decreases. They even experience tracking failures or large drift errors on $L_{10}$ and $L_{11}$.
@@ VII-D3 Results
TABLE VI: Ablation study of line features for mapping. We present the RMSE (m) of our system with and without line features.
EuRoC OIVIO UMA-VI Dark-EuRoC
w/o Lines 0.046 0.072 0.545 0.035
Ours 0.030 0.058 0.441 0.029
VII-D4 Analysis
We find that the VPR methods perform very well on the first type of image but perform poorly on the second type of image. Therefore, their recall rates are very low on P001 and P006, which contain more of the second type of images. This may be because their global features are usually trained on datasets that have a lot of distinct buildings and landmarks, which makes them rely more on such semantic cues to retrieve similar images. By contrast, our system is based on the DBoW method, which only utilizes the low-level local features of images, so it achieves similar performance on the first and second types of images. This also proves the strong generalizability of our system. However, neither our system nor VPR methods can process the pairs with little overlap due to relying on local feature matching. Such image pairs are abundant in the P011.
TABLE VII: Relocalization ablation study. The recall rates (%) of our system with and without the structure graph during map reuse.
@@ VII-E Ablation Study
VII-E1 Line Feature In Mapping
In this section, we evaluate the impact of line features on mapping performance. To this end, we remove line features from our system and compute the RMSE on the EuRoC, OIVIO, UMA-VI, and Dark-EuRoC datasets. The results are summarized in Table VI, where “w/o Lines” refers to our system without line features. As shown in the table, incorporating line features reduces the RMSE by 34.8%, 19.4%, 19.1%, and 17.1% on the four datasets, respectively. These results demonstrate that line features substantially enhance our system and improve the mapping accuracy. We believe this improvement comes from two aspects. On the one hand, a line feature is more likely to be observed across multiple frames compared to a point feature, thereby constraining more frames during pose optimization. This consistent constraint across multiple frames enhances the accuracy of pose estimation. On the other hand, we use the tracking results of a set of points to track a line feature, which makes the tracking of a single line feature more stable than that of a single point.
VII-E2 Structure Graph In Relocalization
We also verify the effectiveness of the proposed relocalization method. This experiment is conducted on the TartanAir Day/Night Localization dataset. We compare the systems with and without the second step proposed in Section VI-B2. The results are presented in Table VII, where “w/o G.” denotes our system without the structure graph, and $N_{\mathcal{C}}$ denotes the candidate number for local feature matching. It shows that using junctions, line features, and structure graphs to filter out relocalization candidates significantly improves recall rates. AirSLAM outperforms w/o G. across all sequences, and when $N_{\mathcal{C}}$ is 3, 5, and 10, the average improvements are 9.6%, 6.0%, and 3.8%, respectively, which demonstrates the effectiveness of the proposed method.
VII-F Efficiency Analysis
Efficiency is essential for robotics applications, so we also evaluate the efficiency of the proposed system. We first compare the running time of our AirSLAM with several SOTA VO and SLAM systems on a computer with an Intel i9-13900 CPU and an NVIDIA RTX 4080 GPU. Then we deploy AirSLAM on an NVIDIA Jetson AGX Orin to verify the efficiency and performance of our system on the embedded platform.
VII-F1 Odometry Efficiency
The VO/VIO efficiency experiment is conducted on the MH_01_easy sequence of the EuRoC dataset. We compare our AirSLAM with several SOTA systems. The loop detection and GBA are disabled from all the systems for a fair comparison. The metrics are the runtime per frame and the CPU usage. The results are presented in Fig. 14a, where 100% CPU usage means 1 CPU core is utilized. It should be additionally noted that DROID-SLAM actually uses 32 CPU cores, and its CPU usage in Fig. 14a is only for a compact presentation. Our system is very efficient, achieving a rate of 73 FPS. In addition, due to extracting and matching features using the GPU, our system requires relatively less CPU resources. We also test the GPU usage. It shows that DROID-SLAM requires about 8GB GPU memory, while our AirSLAM only requires around 3GB.
To investigate the contributions of the CPU and GPU to the efficiency of our system, we measured the total runtime of modules executed on the GPU and CPU using the MH_01_easy sequence and compared it with ORB-SLAM3. The loop detection and GBA are disabled from both systems. The results are presented in Table VIII. It can be seen that our feature detection and matching are highly efficient due to the use of the GPU resources and our system design, which processes only the left image in non-keyframes. Note that we perform the initial pose estimation in the back-end, whereas ORB-SLAM3 includes this process in the front-end. This leads to the differences in the runtime of the “Others” and “Back-end” components between the two systems.
@@ VII-F1 Odometry Efficiency
Figure 14: For odometry efficiency, we disable the loop detection from all the methods. The GPU usage of systems using GPU is also indicated. For mapping efficiency, we compute the total mapping time and compare it with Hloc.
TABLE VIII: Evaluation of CPU and GPU contributions to our system.
Front-end Back-end
FD+FM1 Others
[…]
  • 1
    Total runtime for feature detection and feature matching.   and   represent the modules running on the GPU and CPU, respectively.
@@ VII-F3 Embedded Platform
TABLE IX: Efficiency comparison of our system on two platforms.
@@
We use 8 sequences in the EuRoC dataset to evaluate the efficiency of AirSLAM on the embedded platform. The suffixes, i.e., “-Jetson” and “-PC”, are added to distinguish results on different platforms. On the Jetson, we modify three parameters in our system to improve efficiency. First, we reduced the number of detected keypoints from 350 to 300. Second, we change two parameters in Section V-D to make keyframes sparser, i.e., $\alpha_{1}$ and $\alpha_{2}$ are changed from 0.65 and 0.1 to 0.5 and 0.2, respectively. The other parameters are the same on these two platforms. The comparisons of efficiency and absolute trajectory error (ATE) are presented in Table IX and Fig. 15, respectively. Our AirSLAM can run at a rate of $40\mathrm{Hz}$ on the Jetson while only consuming 2 CPU cores and 989MB GPU memory. We find the runtime of the offline map optimization is very close on these two platforms. This is because AirSLAM-Jetson selects fewer keyframes than AirSLAM-PC, so the loop closure and GBA are faster.
@@ VIII Conclusion
In this work, we presented an efficient and illumination-robust hybrid vSLAM system. To be robust to challenging illumination, the proposed system employs a CNN to detect both keypoints and structural lines. Then these two features are associated and tracked using a GNN. To enhance the efficiency, we proposed PLNet, a unified model capable of simultaneously detecting both point and line features. Furthermore, a multi-stage relocalization method based on both appearance and geometry information was proposed for efficient map reuse. We designed the system with an architecture that includes online mapping, offline optimization, and online relocalization, making it easier to deploy on real robots. Extensive experiments show that the proposed system outperforms other SOTA vSLAM systems in terms of accuracy, efficiency, and robustness in illumination-challenging environments.
Despite its remarkable performance, the proposed system still has limitations. Like other point-line-based SLAM systems, our AirSLAM relies on enough line features, so it is best to apply it to man-made environments. This is because our system was originally designed for warehouse robots. In unstructured environments, the system degrades into a point-only system.